Dublin University Rifle Club

Version Date : April 23, 2002

DURC Competition Attendance Report

Air Rifle Nationals 2002

Distribution

Dr. Trevor West DUCAC Chairman
Aubrey Storey DUCAC Secretary
Terry McAuley DURC President and Director of Sport
Michelle Bennett Physical Recreation Officer
Tony O’Donnell Student Record
Susan Thompson Trinity News

Summary

Venue Wilkinstown Target Shooting Club, Navan, Co.Meath
Date April 13/14, 2002
Discipline 10m Air Rifle

DURC Attendees

Steven Watterson  Mark Dennehy
Ray Reilly  Keith White
Peter Byrne  Nithin Santhosh
Adrian Costigan  Kate Fanning
John Keeney  Rob McAdoo
Victoria McDowell  Doug Woolett
Deirdre Ninaber  Ben Jones

Scores

Steven Watterson 574 578
Mark Dennehy 543 538
Ray Reilly 544 519
Peter Byrne 531 528
Nithin Santhosh 526 524
Adrian Costigan 488 522
Kate Fanning 491 500
John Keeney 506 483
Rob McAdoo 457 422
Victoria McDowell 426 419
Doug Woolett 378 427
Deirdre Ninaber  537
Ben Jones  533

Prizes Won

Steven Watterson Bronze, National Open; 3rd, Class A
Peter Byrne 2nd Class C, National Open; 3rd Class C, National Championships
Nithin Santhosh 3rd ClassC, National Open
Kate Fanning 1st Class D, National Open
John Keeney 2nd Class D, National Open; 3rd Class D, National Championships
Rob McAdoo 3rd Class D, National Open
Deirdre Ninaber 1st Class D, National Championships
Ben Jones 2nd Class C, National Championship

Report

DURC sent a large team of 14 people (13 competitors and one administrative officer) to this year’s Air Rifle National Championships, which were held in Wilkinstown this year for the first time in a decade, in the new NTSA portable range. This range shows how uncomplicated the requirements are for an olympic discipline like 10m Air Rifle. The full range seen here can be disassembled and moved in only three hours, though it does require several people and vans to move the various components. So long as an appropriately sized venue is available, the range can be used. This range should be examined with the new projectile design in mind - it demonstrates how the most important requirement is the physical size of the hall, but also the necessity for permanent fixtures - a three hour setup time would be prohibitive in anything but a temporary range like this one.

The DURC squad shot extremely well, even given the required 5am starts on Saturday and Sunday (necessary to arrive in Wilkinstown with sufficent time to prepare to shoot). All bar Mark Dennehy and Steven Watterson set personal best scores, and Steven held the lead in the National Open until the last few shots. The National Open was held concurrently with the National Championship to allow for international shooters like Steven to compete. Steven managed to secure the bronze medal for DURC in the Open.

The Colours squad used this event as a warm-up and training exercise for the colours match (due to be held in Wilkinstown on April 26th), and given the scores acheived, this was judged to be a resounding success. The squad learnt many valuable lessons in match strategy, and also gained some experience of shooting under pressure. Deirdre Ninaber, shooting in her first ever air rifle match, managed to take first place in class D with a 537, a phenonomally high score for a beginner, and which will see her moved up to class B with her next competition. This bodes very well for the upcoming colours match, which this year is an offically

CUSAI–recognised Intervarsity for the first time.

The DURC Squad. Back Row, left to right: Doug Woolett, Ray Reilly, Peter Byrne, John Keeney, Adrian Costigan, Ben Jones, Mark Dennehy Front Row, left to right: Victoria McDowell, Deirdre Ninaber, Kate Fanning and captain Steven Watterson

Misé le meas,

Mark Dennehy, DURC Secretary